Circular No. 8041

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003A IN UGC 5904 M. Schwartz, Cottage Grove, OR, and W. Li, University of California

at Berkeley, report the LOTOSS (cf. IAUC 7906) discovery of an apparent supernova on an unfiltered image taken under poor conditions with the 0.8-m Tenagra II automated telescope on Jan. 4.5 UT. The new object was confirmed at mag 17.4 on Tenagra (Jan. 5.4) and KAIT (Jan. 3.5) images. SN 2003A is located at α = 10^{h}48^{m}33^{s}.86, δ = +66^{o}22′29′′.4 (equinox 2000.0), which is 0′′.3 east and 6′′.1 north of the nucleus of the northern elliptical component of the interacting galaxy UGC 5904. A KAIT image taken on 2002 Nov. 22.5 showed nothing at this position (limiting mag about 19.0).

SUPERNOVA 2002kf T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center

for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2002kf (cf. IAUC 8040), obtained by M. Calkins on 2002 Dec. 31.34 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova near maximum. Narrow emission lines, from a separate exposure of the nucleus, indicate a recession velocity of 5800 km/s for the host galaxy; using this value, the supernova expansion velocity is ∼ 11100 km/s for Si II (rest 635.5 nm). The spectral-feature age (cf. IAUC 8037) of the supernova is 0 ± 2 days before maximum light.

SUPERNOVA 2002hh IN NGC 6946 P. Chandra and A. Ray, Tata Institute of Fundamental Research,

Mumbai; and S. Bhatnagar, National Radio Astronomy Observatory, Socorro, report the detection of SN 2002hh (cf. IAUC 8005) at 1396.75 MHz (flux density 1.35 ± 0.17 mJy) with the Giant Meterwave Radio Telescope (GMRT) on 2002 Dec. 27. The supernova was not detected on Nov. 30, with the GMRT in the same band (3σ upper limit of 0.42 mJy).

SUPERNOVA 2002jp IN NGC 3313 L. A. G. Monard, Pretoria, S. Africa, reports the following magnitudes

from his unfiltered CCD images: 2002 Oct. 12.10 UT, [18.5; Nov. 12.075, 17.7 (prediscovery); Dec. 2.043, 17.0.

2003 January 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green